# biblio.bib

### v3

@@ 6.3 Approximation of the constant
[…]
@@
[…]
@@ 6.4.1 Number of sources
Intuitively, the most common scenario is that there is a positive entry in $a_{k,k+1}$. In this case the sub-matrix ${(a_{i,j})}_{k\leq i,j\leq n}$ can be re-interpreted as a DOAG with only one source. Indeed, the condition $a_{k,k+1}$ means that upon removing the $(k-1)$ first sources of the DOAG, the decomposition algorithm does not produce any new source, leaving us with a single-source DOAG. We can characterise those matrices: they are made of $(k-1)$ variations of size $(n-k)$ in the first rows, and a size-$(n-k+1)$ labelled transition matrix corresponding to a single source DOAG below them. Any combination of $(k-1)$ such variations and such a matrix can be obtained.
@@
On the other hand we have the matrices such that $a_{k,k+1}=0$. In this case, the $(k-1)$ first rows of the matrix are still variations of size $(n-k)$. The lower part ${(a_{i,j})}_{k\leq i,j\leq n}$ can be seen as a DOAG with at least two sources because its first two columns are empty. Note that here, depending of the $(k-1)$ top variations, we may have restriction on which DOAGs may appear in the lower part. For instance, if $k=2$ and the first row is $(0,0,1,0,0,\ldots,0)$, then the DOAG of size $(n-1)$ without any edge cannot appear in the lower part.
@@ 6.4.2 Number of edges
where $v_{n,p}$ is the number of variations of size $n$ having exactly $p$ zeros, which is equal to $n!/p{!}$ by Lemma 3. Moreover, the tail of their probability distribution is more than exponentially small:
@@
For all integer $p\geq 0$, by inclusion, we have that the number of DOAG labelled transition matrices with exactly $p$ zeros strictly above the diagonal is upper-bounded by the number $\operatorname{\mathrm{VM}}_{n,p}$ of variation matrices with the same property. Moreover, given a size-$(n-1)$ vector $(p_{1},p_{2},\ldots,p_{n-1})$ of non-negative integers such that for all $i$$p_{i}\leq i$, the number of such variation matrices with exactly $p_{n-i}$ zeros in the $i$-th line is
[…]
@@
[…]
@@
[…]
The sum in the last equation is the remainder in the Taylor expansion of order $(q-1)$ of the function $e^{x}$ near zero, evaluated at the point $(n-1)$. By using the integral form of this remainder, we have that
@@
[…]
@@ 7.1 Generating variations
$\mathbb{P}[\text{a uniform variation of size~{}$n$ has~{}$p$ zeros}]=\frac{v_{n,p}}{v_{n}}\propto\frac{\boldsymbol{1}_{\{0\leq p\leq n\}}}{p!}$
by Lemma 3. A possible way to generate a uniform variation is thus to draw a Poisson variable $p$ of parameter $1$ conditioned to be at most $n$, and then shuffling a size $p$ array of zeros concatenated with the identity permutation using the Fisher-Yates algorithm [FY1948]. This is described in Algorithm 3.
@@
4:     $p$leftarrow BoundedPoisson($1,n$)
5:     $A$leftarrow $[0,0,\ldots,0,1,2,\ldots,n-p]$$\triangleright$ array of length $n$, starting with $p$ zeros
6:     for $i=0$ to $n-2$ do
7:         $r$leftarrow Unif($\llbracket i;n-1\rrbracket$)
8:         $A[r]\leftrightarrow A[i]$triangleright Swap entries of indices $r$ and $i$
9:     return $A$
@@
Moreover, when $n$ is large we have $1-p_{n}\sim\tfrac{1}{(n+1)!}$. As a consequence, the expected number of tries of a rejection procedure for sampling conditioned $\text{Poisson}(1)$ variables is $\frac{1}{p_{n}}\leq\frac{e}{2}$. The algorithm described by Knuth in [Knuth1997, page 137] is suitable for our use-case since our Poisson parameter ($1$ here) is small. Furthermore it can be adapted to stop early when values strictly larger than $n$ are found. This is described in Algorithm 4.
@@
5:         $k$leftarrow 00
6:         $p$leftarrow Unif($[0;1]$)
@@
8:              $k$leftarrow $k+1$
[…]
@@
Note that this algorithm relies on real numbers arithmetic. In practice, approximating these numbers by IEEE 754 floating points numbers [ieee754] should introduce an acceptably small error. Indeed, since we only compute products (no sums or subtractions), which generally have few terms, the probability that they introduce an error should not be too far from $2^{-53}$ on a 64-bits architecture. Of course this is only a heuristic argument. A rigorous implementation must keep track of these errors. One possible way would be to use fixed points arithmetic for storing $p$ and to lazily generate the base 2 expansions of the uniform variables at play until we have enough bits to decide how $p$ and $e^{-\lambda}$ compare at line 7. Another way would be to use Ball arithmetic [Hoeven2010, arb] and to increase precision every time the comparison requires more bits. The proofs of correctness and complexity below obviously assume such an implementation.
@@
The correctness of Algorithm 4 follows from the arguments given in [Knuth1997, page 137], which we do not recall here. Regarding Algorithm 3, the for loop at line 6 implements the Fisher-Yates [FY1948] algorithm, which performs a uniform permutation of the contents of the array independently of its contents. In our use-case, this implies that:
@@
$\mathbb{P}[\textsc{BoundedPoisson}(1,n)=p]\cdot\frac{p!}{n!}=\frac{1}{p!\sum_{k=0}^{n}\frac{1}{k!}}\cdot\frac{p!}{n!}=\frac{1}{v_{n}}.\qed$ ∎
@@
Regarding the shuffling happening at line 6 in Algorithm 3, it needs to draw $(n-1)$ uniform integers, respectively smaller or equal to 111122223333, …, $n-1$. At the first order, this incurs a total cost in terms of random bits, of
@@ 7.2 A fast rejection procedure
     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow a zero-filled $n\times n$ matrix
@@
         for $i$ from $1$ to $n-1$ do
              ${(a_{i,j})}_{i<j\leq n}$leftarrow UnifVariation($n-i$)
     until $A$ encodes a DOAG
     return The DOAG corresponding to $A$
@@
As we can see in Theorem 3, the conditions that a variation matrix must satisfy to be a labelled transition matrix concern the shape of the boundary between the zero-filled region between the diagonal and the first positive values above the diagonal. Moreover, we have seen in Theorem 6 that uniform DOAGs tend to have close to $\binom{n}{2}$ edges and thus only a linear number of zeros above the diagonal of their labelled transition matrix. We can thus expect that the area that we have to examine to have access to this boundary should be small. This heuristic argument, hints at a more sparing algorithm that would start by filling the matrix near the diagonal and check its validity early, before generating the content of the whole matrix. This idea, of performing rejection as soon as possible in the generation process, is usually referred to as “anticipated rejection” and also appears in [DFLS2004] and [BPS1994] for instance.
@@
  • for each row of the matrix (i.e. each variation to be sampled), we draw a Poisson variable $p_{i}$ of parameter $1$ and bounded by $(n-i)$;
@@
Figure 15: Ingredient two of the fast rejection-based algorithm: the zeros of the matrix need not be explicitly written. Instead, we interpret the numbers strictly larger than $(n-i-p_{i})$ as zeros. In this example $(n-i)=10$ and $p_{i}=3$ so the numbers 88889999, and $10$ are seen as zeros.
@@
  • for each $2\leq j\leq n$ we need to compute the number $b_{j}=\max\,\{i\ |\ a_{i,j}>0\}$ (or 00 if this set is empty);
@@
A way of implementing this is to start filling each column of the matrix from bottom to top, starting from the column $j=1$ and ending at column $j=n$. For each column, we stop as soon as either a non-zero number is found or the constraints from Theorem 3 are violated. In order to check these constraints, while filling column $j$ from bottom $(i=j-1)$ to top, we halt as soon as either the cell on the left of the current cell, or the current cell is non-zero. The case when the left cell is non-zero corresponds to when $i=b_{j-1}$ and the conditions of Theorem 3 can be checked. Recall that, per the previous point, the zero test in row $i$ is actually $x\mapsto x>n-i-p_{i}$. We shall prove that this process uncovers only a linear number of cells of the matrix, thus allowing to reject invalid matrices in linear expected time. This idea is pictured in Figure 16.
@@
The contents of the ${(p)}_{1\leq i<n}$ vector is also forgotten and each value is to be drawn again before any access. The while loop at line 10 implements the traversal of the matrix described above: at each step, the value of the $a_{i,j}$ is drawn and the conditions of Theorem 3 are checked before proceeding to the next step. The array ${(s_{i})}_{1\leq i\leq n}$ stores the state of each lazy variation generator: $s_{i}$ contains the value of the largest $j$ such that $a_{i,j}$ has been drawn. The second part of the algorithm, starting from line 23, completes the generation of the matrix once its near-diagonal part is known and we know no rejection is possible any more. This includes replacing some values of the matrix by 00 because of ingredient two above.
@@
4:     $A={(a_{i,j})}_{1\leq i,j\leq n}$leftarrow the strictly upper triangular matrix ${(\boldsymbol{1}_{\{j>i\}}\cdot(j-i))}_{1\leq i,j\leq n}$
5:     ${(p_{i})}_{1\leq i<n}$leftarrow uninitialised array
6:     ${(s_{i})}_{1\leq i<n}$leftarrow uninitialised array
7:     $\triangleright$ Anticipated rejection phase
8:         $(i,j)$leftarrow $(1,2)$$\triangleright$ position of the current cell
9:         $p_{1}$leftarrow BoundedPoisson($n-1$)
@@
11:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
@@
17:                  $j$leftarrow $j+1$
18:                  $i$leftarrow $j-1$
19:                  $p_{i}$leftarrow BoundedPoisson($1,n-i$)
@@
27:              $r$leftarrow Unif($\llbracket j;n\rrbracket$)
@@
In the rejection phase, in each column, we draw a certain number of zeros and at most one non-zero value before deciding whether to reject the matrix or to proceed to the next column. As a consequence, when lazily generating a variation matrix, we see at most $(n-1)$ non-zero values and a certain number of zeros that we can trivially upper-bound by the total number of zeros (strictly above the diagonal) in the matrix.
@@ 8 Conclusion and perspectives
So far, we have only approached DOAGs via enumerative tools and ad. hoc. asymptotic techniques. A common and powerful tool in combinatorics is the use of generating function to tackle not only asymptotic estimation, but also the convergence in distribution of parameters, such as the number of edges. The book [FS2009] is a reference on this topic. Classical approaches using ordinary, exponential, or even graphical ([PD2019]) generating functions fail in our context due to the super factorial behaviour of the number of DOAGs. It remains an open question whether it is possible to design a generating function approach to our objects, which could help obtaining finer estimates, not only over $D_{n}$, but also over the low of the number of edges.
@@
An interesting question that is left open by our work is the case of the multi-graph variant of this model: what happens if multiple edges are allowed between two given vertices? This makes the analysis more challenging since there is now an infinite number of objects with $n$ vertices. We thus must change our point of view and take the number $m$ of edges into account in addition to, or instead of, the number of vertices. Estimating the number and behaviour of DOAGs as well as their multi-graph counterpart, when both parameters $n$ and $m$ grow remains an open question and will certainly yield very different results depending on how $n$ and $m$ grow in relation to each other. We argue that the model of multi-edge DOAGs is natural, maybe even more so than that of DOAGs, since they encode (partially) compacted plane trees. Quantitative aspects of tree compaction, in particular the typical compression rate, has been studied in the past [FSS1990] in a general setting. However, the dual point of view that consists in studying already compacted structures directly is a more recent topic, see [EFW2021] and [GW2024] for instance. The class of multi-edge DOAGs generalises the classes studied in those two papers. Moreover, being able to sample them efficiently would give a tool to reach every possible case (including those with double edges) when testing programs manipulating compacted trees (such as compilers) via random generation.
Another interesting question is that of the connectivity. We do not provide a way to count connected DOAGs directly here. However we have already proved that, in the uniform model, they are connected with high probability since they have only one source with high probability. Moreover, since $D_{n}$ grows extremely fast, we can also foresee that a uniform DOAG of size $n$ with two connected components will typically have one tiny component of size $1$ and a big component of size $n-1$, and that the asymptotic estimations of such graphs is straightforward. This implies that sampling a uniform connected DOAG with $n$ vertices is already possible, and efficient, by rejection and the question of their direct enumeration is thus mostly of mathematical interest.
@@
Finally, it is also natural to wonder whether our successful approach at efficiently sampling DOAG applies to labelled DAGs. Of course, the asymptotics of DAGs is known [Robinson1973]. But if a matrix encoding similar to ours is feasible, that is an encoding whose combinatorial properties are understood well enough to avoid introduction any bias, then it might be possible to devise an efficient, pre-computation-free, uniform sampler for DAGs.
@@
Notable is that the approach presented in [KM2015] can also be seen as a way to work with matrices while maintaining uniformity by using a combinatorial encoding using ordered integer partitions. A caveat however is that their approach still requires a costly pre-processing, which we seek to avoid using a rejection-based approach.
[…]

### v4

@@ 6.3 Approximation of the constant
Note that the $D_{n}^{\star}$ sequence from Theorem 4 corresponds to $D_{n,1}$ and that $D_{n}=\sum_{k=1}^{n}\hskip-1.0ptD_{n,k}$. Using the numbers computed by this algorithm, we plotted the first $250$ values of the sequences $D_{n}$ and $D_{n}^{\star}$ normalised by $n^{-1/2}e^{n-1}\text{¡}{(n-1)}!$ which shows the convergence to the constant $c$ from Theorem 4. We also note that the convergence looks faster for the sequence $D_{n}^{\star}$. This suggests that the constant can be approximated by $c\approx 0.4967$. Figure 12 shows this plot as well as a zoomed-in version near $\frac{1}{2}$ for $n\geq 200$.
@@
Figure 12: The first values of the sequences ${\frac{D_{n}\sqrt{n}}{\text{¡}{(n-1)}!e^{n-1}}}$ in blue and ${\frac{D_{n}^{\star}\sqrt{n}}{\text{¡}{(n-1)}!e^{n-1}}}$ in red.
@@ 6.4.1 Number of sources
Intuitively, the most common scenario is that there is a positive entry in $a_{k,k+1}$. In this case the sub-matrix ${(a_{i,j})}_{k\leq i,j\leq n}$ can be re-interpreted as a DOAG with only one source. Indeed, the condition $a_{k,k+1}$ means that upon removing the (k1)(k-1) first sources of the DOAG, the decomposition algorithm does not produce any new source, leaving us with a single-source DOAG. We can characterise those matrices: they are made of (k1)(k-1) variations of size (nk)(n-k) in the first rows, and a size-(nk+1)(n-k+1) labelled transition matrix corresponding to a single source DOAG below them. Any combination of (k1)(k-1) such variations and such a matrix can be obtained.
@@
On the other hand we have the matrices such that $a_{k,k+1}=0$. In this case, the (k1)(k-1) first rows of the matrix are still variations of size (nk)(n-k). The lower part ${(a_{i,j})}_{k\leq i,j\leq n}$ can be seen as a DOAG with at least two sources because its first two columns are empty. Note that here, depending of the (k1)(k-1) top variations, we may have restriction on which DOAGs may appear in the lower part. For instance, if $k=2$ and the first row is $(0,0,1,0,0,\ldots,0)$, then the DOAG of size (n1)(n-1) without any edge cannot appear in the lower part.
@@ 6.4.2 Number of edges
where $v_{n,p}$ is the number of variations of size $n$ having exactly $p$ zeros, which is equal to n!/p!n!/p{!} by Lemma 3. Moreover, the tail of their probability distribution is more than exponentially small:
@@
For all integer $p\geq 0$, by inclusion, we have that the number of DOAG labelled transition matrices with exactly $p$ zeros strictly above the diagonal is upper-bounded by the number $\operatorname{\mathrm{VM}}_{n,p}$ of variation matrices with the same property. Moreover, given a vector $(p_{1},p_{2},\ldots,p_{n-1})$ of non-negative integers such that for all $ip_{i}\leq i$, the number of such variation matrices with exactly $p_{n-i}$ zeros in the $i$-th line is
$\prod_{i=1}^{n-1}v_{i,p_{i}}=\prod_{i=1}^{n-1}\frac{i!}{p_{i}!}=\text{¡}{(n-1)}!\prod_{i=1}^{n-1}\frac{1}{p_{i}!}.$
@@
$\operatorname{\mathrm{VM}}_{n,p}=\text{¡}{(n-1)}!\sum_{\begin{subarray}{c}p_{1}+p_{2}+\cdots+p_{n-1}=p\\ \text{for all~$i$},0\leq p_{i}\leq i\end{subarray}}\prod_{i=1}^{n-1}\frac{1}{p_{i}!}\leq\text{¡}{(n-1)}!\sum_{\begin{subarray}{c}p_{1}+p_{2}+\cdots+p_{n-1}=p\\ \text{for all~$i$},0\leq p_{i}\end{subarray}}\prod_{i=1}^{n-1}\frac{1}{p_{i}!}\cdot$
@@
$\mathbb{P}_{n}[\text{a uniform DOAG has at most~$\binom{n}{2}-q$ zeros}]\leq\frac{\text{¡}{(n-1)}!}{D_{n}}\sum_{p\geq q}\frac{{(n-1)}^{p}}{p!}\cdot$ (15)
The sum in the last equation is the remainder in the Taylor expansion of order (q1)(q-1) of the function $e^{x}$ near zero, evaluated at the point (n1)(n-1). By using the integral form of this remainder, we have that
@@
$\mathbb{P}_{n}[\text{a uniform DOAG has at most~$\binom{n}{2}-\lambda(n-1)$ zeros}]$ $\leq\frac{\text{¡}{(n-1)}!e^{n-1}}{D_{n}}\sum_{p\geq q}\frac{{(n-1)}^{q}}{q!}$
@@ 7.1 Generating variations
[…]
by Lemma 3. A possible way to generate a uniform variation is thus to draw a Poisson variable $p$ of parameter $1$ conditioned to be at most $n$, and then shuffling a size $p$ array of zeros concatenated with the identity permutation using the Fisher-Yates algorithm [14]. This is described in Algorithm 3.
@@
4:  $p$ $\leftarrow$ BoundedPoisson($1,n$)
5:  AA $\leftarrow$ $0,0,\ldots,0,1,2,\ldots,n-p]\triangleright$ array of length $n$, starting with $p$ zeros
6:  for $i=0$ to n2n-2 do
7:   $r$ $\leftarrow$ Unif($\llbracket i;n-1\rrbracket$)
8:   $A[r]\leftrightarrow A[i]$ $\triangleright$ Swap entries of indices $r$ and $i$
9:  return AA
@@
Moreover, when $n$ is large we have $1-p_{n}\sim\tfrac{1}{(n+1)!}$. As a consequence, the expected number of tries of a rejection procedure for sampling conditioned $\text{Poisson}(1)$ variables is $\frac{1}{p_{n}}\leq\frac{e}{2}$. The algorithm described by Knuth in [28, page 137] is suitable for our use-case since our Poisson parameter ($1$ here) is small. Furthermore it can be adapted to stop early when values strictly larger than $n$ are found. This is described in Algorithm 4.
@@
5:   $k$ $\leftarrow$ 0
6:   $p$ $\leftarrow$ Unif($[0;1]$)
@@
8:     $k$ $\leftarrow$ $k+1$
9:     $p$ $\leftarrow$ $p\cdot\text{{Unif}($[0;1]$)}$
@@
Note that this algorithm relies on real numbers arithmetic. In practice, approximating these numbers by IEEE 754 floating points numbers [40] should introduce an acceptably small error. Indeed, since we only compute products (no sums or subtractions), which generally have few terms, the probability that they introduce an error should not be too far from $2^{-53}$ on a 64-bits architecture. Of course this is only a heuristic argument. A rigorous implementation must keep track of these errors. One possible way would be to use fixed points arithmetic for storing $p$ and to lazily generate the base 2 expansions of the uniform variables at play until we have enough bits to decide how $p$ and $e^{-\lambda}$ compare at line 7. Another way would be to use Ball arithmetic [24, 26] and to increase precision every time the comparison requires more bits. The proofs of correctness and complexity below obviously assume such an implementation.
@@
The correctness of Algorithm 4 follows from the arguments given in [28, page 137], which we do not recall here. Regarding Algorithm 3, the for loop at line 6 implements the Fisher-Yates [14] algorithm, which performs a uniform permutation of the contents of the array independently of its contents. In our use-case, this implies that:
@@
$\mathbb{P}[\textsc{BoundedPoisson}(1,n)=p]\cdot\frac{p!}{n!}=\frac{1}{p!\sum_{k=0}^{n}\frac{1}{k!}}\cdot\frac{p!}{n!}=\frac{1}{v_{n}}.\qed$
@@
Regarding the shuffling happening at line 6 in Algorithm 3, it needs to draw (n1)(n-1) uniform integers, respectively smaller or equal to 112233, …, n1n-1. At the first order, this incurs a total cost in terms of random bits, of
@@ 7.2 A fast rejection procedure
  $A={(a_{i,j})}_{1\leq i,j\leq n}$ $\leftarrow$ a zero-filled $n\times n$ matrix
@@
   for $i$ from $1$ to n1n-1 do
     ${(a_{i,j})}_{i<j\leq n}$ $\leftarrow$ UnifVariation(nin-i)
  until AA encodes a DOAG
  return The DOAG corresponding to AA
@@
As we can see in Theorem 3, the conditions that a variation matrix must satisfy to be a labelled transition matrix concern the shape of the boundary between the zero-filled region between the diagonal and the first positive values above the diagonal. Moreover, we have seen in Theorem 6 that uniform DOAGs tend to have close to $\binom{n}{2}$ edges and thus only a linear number of zeros above the diagonal of their labelled transition matrix. We can thus expect that the area that we have to examine to have access to this boundary should be small. This heuristic argument, hints at a more sparing algorithm that would start by filling the matrix near the diagonal and check its validity early, before generating the content of the whole matrix. This idea, of performing rejection as soon as possible in the generation process, is usually referred to as “anticipated rejection” and also appears in [10] and [1] for instance.
@@
  • for each row of the matrix (i.e. each variation to be sampled), we draw a Poisson variable $p_{i}$ of parameter $1$ and bounded by (ni)(n-i);
@@
Figure 15: Ingredient two of the fast rejection-based algorithm: the zeros of the matrix need not be explicitly written. Instead, we interpret the numbers strictly larger than $(n-i-p_{i})$ as zeros. In this example (ni)=10(n-i)=10 and $p_{i}=3$ so the numbers 8899, and $10$ are seen as zeros.
@@
  • for each $2\leq j\leq n$ we need to compute the number $b_{j}=\max\,\{i\ |\ a_{i,j}>0\}$ (or 0 if this set is empty);
@@
A way of implementing this is to start filling each column of the matrix from bottom to top, starting from the column $j=1$ and ending at column $j=n$. For each column, we stop as soon as either a non-zero number is found or the constraints from Theorem 3 are violated. In order to check these constraints, while filling column $j$ from bottom (i=j1)(i=j-1) to top, we halt as soon as either the cell on the left of the current cell, or the current cell is non-zero. The case when the left cell is non-zero corresponds to when $i=b_{j-1}$ and the conditions of Theorem 3 can be checked. Recall that, per the previous point, the zero test in row $i$ is actually $x\mapsto x>n-i-p_{i}$. We shall prove that this process uncovers only a linear number of cells of the matrix, thus allowing to reject invalid matrices in linear expected time. This idea is pictured in Figure 16.
@@
The contents of the ${(p)}_{1\leq i<n}$ vector is also forgotten and each value is to be drawn again before any access. The while loop at line 10 implements the traversal of the matrix described above: at each step, the value of the $a_{i,j}$ is drawn and the conditions of Theorem 3 are checked before proceeding to the next step. The array ${(s_{i})}_{1\leq i\leq n}$ stores the state of each lazy variation generator: $s_{i}$ contains the value of the largest $j$ such that $a_{i,j}$ has been drawn. The second part of the algorithm, starting from line 23, completes the generation of the matrix once its near-diagonal part is known and we know no rejection is possible any more. This includes replacing some values of the matrix by 0 because of ingredient two above.
@@
4:  $A={(a_{i,j})}_{1\leq i,j\leq n}$ $\leftarrow$ the strictly upper triangular matrix ${(\boldsymbol{1}_{\{j>i\}}\cdot(j-i))}_{1\leq i,j\leq n}$
5:  ${(p_{i})}_{1\leq i<n}$ $\leftarrow$ uninitialised array
6:  ${(s_{i})}_{1\leq i<n}$ $\leftarrow$ uninitialised array
[…]
8:   $(i,j)$ $\leftarrow$ $(1,2)\triangleright$ position of the current cell
9:   $p_{1}$ $\leftarrow$ BoundedPoisson(n1n-1)
@@
11:     $r$ $\leftarrow$ Unif($\llbracket j;n\rrbracket$)
@@
17:      $j$ $\leftarrow$ $j+1$
18:      $i$ $\leftarrow$ j1j-1
19:      $p_{i}$ $\leftarrow$ BoundedPoisson(1,ni1,n-i)
@@
27:     $r$ $\leftarrow$ Unif($\llbracket j;n\rrbracket$)
@@
In the rejection phase, in each column, we draw a certain number of zeros and at most one non-zero value before deciding whether to reject the matrix or to proceed to the next column. As a consequence, when lazily generating a variation matrix, we see at most (n1)(n-1) non-zero values and a certain number of zeros that we can trivially upper-bound by the total number of zeros (strictly above the diagonal) in the matrix.
@@ 8 Conclusion and perspectives
So far, we have only approached DOAGs via enumerative tools and ad. hoc. asymptotic techniques. A common and powerful tool in combinatorics is the use of generating function to tackle not only asymptotic estimation, but also the convergence in distribution of parameters, such as the number of edges. The book [15] is a reference on this topic. Classical approaches using ordinary, exponential, or even graphical ([9]) generating functions fail in our context due to the super factorial behaviour of the number of DOAGs. It remains an open question whether it is possible to design a generating function approach to our objects, which could help obtaining finer estimates, not only over $D_{n}$, but also over the low of the number of edges.
@@
An interesting question that is left open by our work is the case of the multi-graph variant of this model: what happens if multiple edges are allowed between two given vertices? This makes the analysis more challenging since there is now an infinite number of objects with $n$ vertices. We thus must change our point of view and take the number $m$ of edges into account in addition to, or instead of, the number of vertices. Estimating the number and behaviour of DOAGs as well as their multi-graph counterpart, when both parameters $n$ and $m$ grow remains an open question and will certainly yield very different results depending on how $n$ and $m$ grow in relation to each other. We argue that the model of multi-edge DOAGs is natural, maybe even more so than that of DOAGs, since they encode (partially) compacted plane trees. Quantitative aspects of tree compaction, in particular the typical compression rate, has been studied in the past [16] in a general setting. However, the dual point of view that consists in studying already compacted structures directly is a more recent topic, see [12] and [21] for instance. The class of multi-edge DOAGs generalises the classes studied in those two papers. Moreover, being able to sample them efficiently would give a tool to reach every possible case (including those with double edges) when testing programs manipulating compacted trees (such as compilers) via random generation.
Another interesting question is that of the connectivity. We do not provide a way to count connected DOAGs directly here. However we have already proved that, in the uniform model, they are connected with high probability since they have only one source with high probability. Moreover, since $D_{n}$ grows extremely fast, we can also foresee that a uniform DOAG of size $n$ with two connected components will typically have one tiny component of size $1$ and a big component of size n1n-1, and that the asymptotic estimations of such graphs is straightforward. This implies that sampling a uniform connected DOAG with $n$ vertices is already possible, and efficient, by rejection and the question of their direct enumeration is thus mostly of mathematical interest.
@@
Finally, it is also natural to wonder whether our successful approach at efficiently sampling DOAG applies to labelled DAGs. Of course, the asymptotics of DAGs is known [36]. But if a matrix encoding similar to ours is feasible, that is an encoding whose combinatorial properties are understood well enough to avoid introduction any bias, then it might be possible to devise an efficient, pre-computation-free, uniform sampler for DAGs.
@@
Notable is that the approach presented in [31] can also be seen as a way to work with matrices while maintaining uniformity by using a combinatorial encoding using ordered integer partitions. A caveat however is that their approach still requires a costly pre-processing, which we seek to avoid using a rejection-based approach.
Acknowledgements
We would like to thank one of the anonymous referees for pointing us at related work that we had missed in the field of Statistics and Bayesian inference, and appeared to have a strong connection with our line of work.
This research received financial support from the LIA IFUMI and the ANR PPS project ANR-19-CE48-0014.